OpenStudy (anonymous):

Log Help last one 2e^(x-3)+7=27

OpenStudy (anonymous):

\[ 2e^{x-3}+7=27\\ 2e^{x-3}=20\\ e^{x-3}=10\\ \ln \ e^{x-3}=\ln10\\ x=3 +\ln(10)\\ 5.3026 \]